OpenStudy (anonymous):

20. What affect does h = 4 have on the parent function y = 1/x A. Graph moves 4 units up B. Vertical stretch by a factor of |4| C. Graph moves 4 units to the right D. Graph moves 4 units left

OpenStudy (anonymous):

Where is 'h?'

OpenStudy (anonymous):

"What affect does h = 4 have on the parent function y = 1/x" Where can I find this 'h' that you're asking about?

OpenStudy (anonymous):

How long have you been studying this topic? Are you sure you haven't seen other formulas that have an 'h' in them?

OpenStudy (anonymous):

There is a standard formula that I know \[\large y=\frac{a}{x-h}+k\] but I want to make sure that's the same one that applies here.

OpenStudy (anonymous):

y=1/x is what you get if a=1, and h=k=0.

OpenStudy (anonymous):

Ah, ok, well that form is pretty standard in mathematics, so I guess it's the right one. In that case, an h=4 causes a horizontal shift. The best way for you to see this is to graph both y=1/x and y=1/(x-4) and see what the transformation looks like.

OpenStudy (anonymous):

Yes. It is similar to a quadratic in vertex form y=a(x-h)^2+k where 'h' is also a horizontal translation.

OpenStudy (anonymous):

Should post this as a new question . . .

OpenStudy (anonymous):

Ok, so look again at the standard equation y=a/(x-h)+k All else being equal, and only changing y=1/x to y=3/x, what do you suppose will happen to y?

OpenStudy (anonymous):

Maybe think of it like this: y=f(x)=1/x 3/x = 3*f(x).

OpenStudy (anonymous):

Right, because all the y's are being multiplied by 3.

OpenStudy (anonymous):

(and y is the vertical direction, yes?)

OpenStudy (anonymous):

Make sense?

OpenStudy (anonymous):

Wonderful! Thank you. While we're on the subject, what do you suppose happens if you change 'k'?

OpenStudy (anonymous):

Use the same reasoning as before. Assume everything else is the same, but now you have y=1/x +k... The 1/x part stays the same, so the change is y=(same as before)+k.

OpenStudy (anonymous):

Yes, but in which direction?

OpenStudy (anonymous):

What is being added to, x or y?